A particle completes 1.5 revolution in a circular path of radius 2 centimetre the angular displacement of particle will be (in radian)

Grade:12

3 Answers

jitender
6 years ago
Angular displacement , ¤ = 2£NWhere £ = 3.14 = 22/7 = pieAnd N= no. Of revolutions.Now N=1.5 =3/2¤=2£3/2=3£ = 9.42 rad
Arun
6 years ago
Dear Shubham
 
As angular displacement = \theta = angle moved
1.5 revolutions = 1 full revolution and 1/2 revolution
But 1 full revolution = 2 \large \pi readian
 
Hence total angle covered = 2\large \pi\large \pi = 3\large \pi radian
 
Hence angular displacement = 3 \large \pi radian
 
Regards
Arun (askIITians forum expert)
 
Abhay
5 years ago
It is very simple..
It can be done by unitary method!
We know, 
1 revolution  =   2π radians
So,  10 radian  =   1/2π×10
                          =     5/π Radian
 
HOPE YOU GOT THE EASIEST WAY !!!! 👍